I need some help on this problem:
I need to write 58% as a fraction or a mixed number.
Thanks, Ashley

Found 2 solutions by fractalier, tutorcecilia:
Answer by fractalier(6550)

58% = 58/100 = 29/50

Answer by tutorcecilia(2152)

58% = 58/100 = 29/50